OpenStudy (anonymous):

A astronaut has 100 kg of mass. Unfortunately the mass of his body was spread out in a sphere. After that, the density inside the sphere is 2x10^-28 kg/m^3. What is the radius of the sphere where the mass of the astronaut was spread out?

OpenStudy (anonymous):

Density = MASS/VOLUME

OpenStudy (anonymous):

use it to find the volume of the sphere

OpenStudy (anonymous):

I just set up the following relation 6x10^-28=100kg/V

OpenStudy (anonymous):

Yes u r correct. PLZ CONTINUE

OpenStudy (anonymous):

So, with some algebric manipulation i found this relation V=1x10^2/6x10^-28 and i got V=1/6x10^30.

OpenStudy (anonymous):

I think this is the volume of the sphere.

OpenStudy (anonymous):

Now, Volume of a sphere=4/3 pi r^3

OpenStudy (anonymous):

Use it to find r

OpenStudy (anonymous):

I just plug it in the formual, right?

OpenStudy (anonymous):

YEp, 1/6x10^30 = 4/3 pi r^3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Just one moment.

OpenStudy (anonymous):

i got 1,46459

OpenStudy (anonymous):

But I got different

OpenStudy (anonymous):

What do you got?

OpenStudy (anonymous):

i write the equation in this order. cubic root of 1/6x10^30=4/3phi*r

OpenStudy (anonymous):

Now i got 4,12740906x10^9 m

OpenStudy (anonymous):

Density= MASS/ VOLUME 2x10^-28=100/VOLUME VOLUME=50 * 10 ^28 4/3 pi r^3 = 50 * 10^28 r^3=4923725109

OpenStudy (anonymous):

That ´s right.

OpenStudy (anonymous):

r=4923725109

OpenStudy (anonymous):

How do u calculate the 2x10^-28=100/VOLUME. Plz do it step by step.

OpenStudy (anonymous):

DEnsity =MASS/ VOLUME

OpenStudy (anonymous):

Now, substitute the given values

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)yw

OpenStudy (anonymous):

But sorry i can´t understant how do you calculate 2x10^-28=100/VOLUME. May you do it plz?

OpenStudy (anonymous):

When i calculate V i have found 50/3x10^28. How did you got 50 * 10 ^28. PLz.

OpenStudy (anonymous):

100/2 =50 right?

OpenStudy (anonymous):

Yes, but where did you find the number two? If we have the division 100/6?

OpenStudy (anonymous):

Did you divide 6/3. I got it. now.